OpenStudy (anonymous):

The average rate of change of the function between x = 10 to x = 16 is ______ m/s and represents the person's speed. time (second) x 10 13 16 19 distance (meters) f(x) 60 78 96 114

OpenStudy (anonymous):

Slope formula. (10,60) and (16,96)

OpenStudy (anonymous):

f(x) is just y, So if you have listed them in the correct order, 10 seconds (x) would be 60 meters (y)

OpenStudy (anonymous):

slope form ? so i would just do 96-60 = 36 over 16 -10 =6?

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

So yes, \[\frac{ 96-60 }{ 16-10 }\]

OpenStudy (anonymous):

36/6 = 6 Is that the answer?

OpenStudy (anonymous):

Yes, the slope would be 6 or 6/1 ( same thing)

OpenStudy (anonymous):

so would 6 be the average rate of change between the two?

OpenStudy (anonymous):

Average Rate of change is basically just slope, those time and distance you posted can be looked at as (x,y) coordinates. Yes 6 would be the rate of change.

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (anonymous):

Np=)